# Derail Yourself: Multi-turn LLM Jailbreak Attack through self-discovered clues

### v1

@@ Abstract
This study exposes the safety vulnerabilities of Large Language Models (LLMs) in multi-turn interactions, where malicious users can obscure harmful intents across several queries. We introduce ActorAttack, a novel multi-turn attack method inspired by actor-network theory, which models a network of semantically linked actors as attack clues to generate diverse and effective attack paths toward harmful targets. ActorAttack addresses two main challenges in multi-turn attacks: (1) concealing harmful intents by creating an innocuous conversation topic about the actor, and (2) uncovering diverse attack paths towards the same harmful target by leveraging LLMs’ knowledge to specify the correlated actors as various attack clues. In this way, ActorAttack outperforms existing single-turn and multi-turn attack methods across advanced aligned LLMs, even for GPT-o1. We will publish a dataset called SafeMTData, which includes multi-turn adversarial prompts and safety alignment data, generated by ActorAttack. We demonstrate that models safety-tuned using our safety dataset are more robust to multi-turn attacks. Code is available at https://github.com/renqibing/ActorAttack.
WARNING: This paper contains unsafe model responses.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities, but they can be misused for both benefit and harm, like social engineering, developing biological weapons, and cyberattacks (Bommasani et al., 2021; Weidinger et al., 2022). To thoroughly investigate the safety vulnerabilities of LLMs, it is critical to discover diverse attack ways that can elicit harmful or inappropriate responses. Current attack methods are mainly single-turn based, which elicit harmful responses from the victim LLM within one turn of the conversation (Wei et al., 2024; Chao et al., 2023; Zeng et al., 2024; Zou et al., 2023b). However, in real-world scenarios, interactions between users and LLMs typically unfold over multiple conversational turns (ShareGPT, 2023).
@@
Figure 1: (a): A real-world example of a multi-turn attack generated by our method compared with the single-turn attack baseline. (b) & (c): Schematic comparison between our method and another multi-turn attack baseline. Each triangle box represents an attack clue, which describes some object related to the harmful target, as a hint for a multi-turn attack. The series of white circles represent a sequence of thoughts about how to finish our multi-turn attack step by step. See concrete examples of how to construct the network, and how to infer the attack chain in Fig. 2 and Fig. 3.
Identifying and dealing with the potential harms and misuse risks in multi-turn interactions is an open research question. Unlike single-turn attacks, where the malicious intent is clear in the prompt, multi-turn interactions enable the malicious users to hide their intentions. For example, as shown in Fig. 1 (a), the user starts with a neutral query like “Who is Ted Kaczynski?” (a terrorist who has bomb-making activities). In each follow-up question, the user induces the victim model to provide more harmful details based on its previous response. Although all the follow-up questions are still innocuous, the user finally obtains the knowledge of bomb-making.
The above example reveals the safety risks in multi-turn conversations, while there are two main challenges in designing such attacks. First, attackers need to hide harmful intent to avoid detection. Second, multi-turn conversations give attackers more opportunities to act, allowing multiple possible attack paths for the same target. The challenge is how to discover these paths to reveal additional safety vulnerabilities. To resolve the first challenge, as shown in Fig. 1 (c), Crescendo implements its attack by gradually guiding benign initial queries towards more harmful topics, based on the fixed and human-crafted seed instances (Russinovich et al., 2024). The performance of Crescendo depends on the quality and relevance of the seed instances with the test cases. If the test cases differ from the seed examples, Crescendo may not generate effective attacks well. Moreover, Crescendo generates different attack paths via random trials, but these paths tend to be biased toward the seed instances and lack diversity, thus not effectively addressing the second challenge.
In this paper, we propose an effective and diverse multi-turn attack method, called ActorAttack. Inspired by Latour’s actor-network theory (Latour, 1987), we explicitly model a network where each node (actor) is semantically linked with the harmful target (e.g., the actor, Ted Kaczynski, who builds bombs for terrorism, is correlated with the target of building a bomb.). These actors and their relationships with the harmful target constitute our attack clues, and we hide the harmful intent in the innocuous multi-turn conversation about the actor. Notably, as illustrated in Fig. 1 (b), we propose automating the discovery of attack clues by leveraging the knowledge of LLMs. Selecting an attack clue, ActorAttack then infers the attack chain, which describes how to achieve harmful targets step by step (Fig. 3). Following the attack chain, ActorAttack generates queries, which can guide LLMs’ responses to become increasingly relevant to the harmful target until reaching it.
Overall, our network design helps improve the diversity of our attack on two levels: (1) inter-network diversity: our attacker model generates target-specific networks for various harmful targets; (2) intra-network diversity: Inside the network, we categorize six distinct types of nodes (actors) based on their relationship to the harmful target and each type of nodes leads to different attack paths (Fig. 2). Experimental results show that ActorAttack finds higher-quality attacks from more diverse attack paths, and is effective over both single-turn and multi-turn attack baselines across various aligned LLMs, even for GPT-o1 (OpenAI, 2024b) whose advanced reasoning improves safety. We find that though GPT-o1 identifies our harmful intent and shows it should follow the safety policies in its chain of thought, it still outputs unsafe content, revealing the potential conflict between its helpfulness and safety goals against our attack.
Finally, we construct a dataset SafeMTData, which includes both multi-turn adversarial prompts and safety alignment data, generated by ActorAttack, as a complementary to the existing single-turn safety alignment datasets (Ji et al., 2024; Bai et al., 2022). We find that performing safety fine-tuning on our safety dataset greatly improves the robustness of LLMs against both ActorAttack and Crescendo, while there exists a trade-off between utility and safety.  [^1] Our data can be found at https://huggingface.co/datasets/SafeMTData/SafeMTData.
@@ 2 Related Work
Multi-turn Attacks. Multi-turn attacks are less covered in the literature, though there have been several works to reveal the safety risks in the multi-turn dialogue scenario. One multi-turn attack strategy is the fine-grained task decomposition, which decomposes the original malicious query into several less harmful sub-questions (Yu et al., 2024; Zhou et al., 2024; Liu et al., 2024d). While this decomposition strategy successfully circumvents current safety mechanisms, it may be easily mitigated by including these finer-grained harmful queries in safety training data. Alternatively, researchers propose to use human red teamers to expose vulnerabilities of LLMs against multi-turn attacks (Li et al., 2024b). Moreover,  Yang et al. (2024) depends on the heuristics from  (Chao et al., 2024) and its seed examples to implement its attacks. The most relevant to our work is Crescendo (Russinovich et al., 2024), which gradually steers benign initial queries towards more harmful topics. The implementation of Crescendo is based on the fixed and human-crafted seed instances, making it challenging to generate diverse and effective attacks (Section 4.3, Fig. 5). By contrast, we propose to discover diverse attack clues inside the model’s prior knowledge. We further model the attack clues via a network and classify these clues into different types, bringing a greater coverage of possible attack paths. Moreover, the inherent semantic correlation between our attack clues and our attack target ensures effectiveness.
Defenses for LLMs. To ensure LLMs safely follow human intents, various defense measures have been developed, including prompt engineering (Xie et al., 2023; Zheng et al., 2024), aligning models with human values (Ouyang et al., 2022; Bai et al., 2022; Rafailov et al., 2024; Meng et al., 2024; Yuan et al., 2024a), model unlearning (Li et al., 2024c; Zhang et al., 2024b), representation engineering(Zou et al., 2024a) and implementing input and output guardrails (Dubey et al., 2024; Inan et al., 2023; Zou et al., 2024b). Specifically, input and output guardrails involve input perturbation (Robey et al., 2023; Cao et al., 2023; Liu et al., 2024e), safety decoding (Xu et al., 2024), and jailbreak detection (Zhang et al., 2024c; Yuan et al., 2024c; Phute et al., 2023; Alon & Kamfonas, 2023; Jain et al., 2023; Hu et al., 2024). Priority training also shows its effectiveness by training LLMs to prioritize safe instructions (Lu et al., 2024; Wallace et al., 2024; Zhang et al., 2023).
3 Method: generate multi-turn attack through self-discovered clues
Overview. We propose a two-stage approach to automatically find attack clues and generate multi-turn attacks. The first stage consists of network construction around the harmful target, where every network node can be used as an attack clue (Fig. 2). The second stage includes the attack chain generation based on the attack clue and the multi-turn query generation (Fig. 3).
@@ 3 Method: generate multi-turn attack through self-discovered clues
Figure 2: Druing the pre-attack stage, ActorAttack first leverages the knowledge of LLMs to instantiate our conceptual network ${\mathcal{G}}_{concept}$ as ${\mathcal{G}}_{inst}$ as a two-layer tree. The leaf nodes of ${\mathcal{G}}_{inst}$ are specific actor names. ActorAttack then samples actors and their relationships with the harmful target as our attack clues.
3.1 Pre-attack: find attack clues
Inspired by Latour’s actor-network theory, we propose a conceptual network ${\mathcal{G}}_{concept}$ to categorize various types of actors correlated with the harmful target. These actors can be exploited as our attack clues, and we leverage the knowledge of LLMs to specify these clues. Latour (1987) claim that everything exists in networks of relationships, and is influenced by human and non-human actors in the network. For better coverage of possible attack clues, as illustrated in Fig. 2, we identify six types of actors in terms of how they influence the harmful target in a network, e.g., Distribution corresponds to actors who spread harmful behaviors or information across the network. Inside each type of actor, we further consider both human and non-human actors, e.g., human actors can include historical figures and influential people and non-human actors can be books, media, or social movements.
To be specific, we model our network into a two-layer tree, whose root node is the harmful target $x$, and the nodes in the first layer represent the six abstract types. The leaf nodes are specific actor names. Each leaf node $v_{i}$ and the edge between the leaf node and the parent node depict the relationship between the actor and the harmful target, which constitutes an attack clue $c_{i}$. Since LLMs are trained on vast amounts of text data, we view LLMs as “knowledge base” to instantiate our conceptual network ${\mathcal{G}}_{concept}$ as ${\mathcal{G}}_{inst}$, i.e., ${\mathcal{G}}_{inst}\sim p(x,{\mathcal{G}}_{concept};\theta)$. Finally, we extract our diverse attack clue set $C$=$[c_{1},\ldots,c_{n}]$ from ${\mathcal{G}}_{inst}$, i.e., $C\sim{\mathcal{G}}_{inst}$.
3.2 In-attack: first reason then attack
Based on the identified attack clue, we perform our multi-turn attacks in three steps. The first step is about inferring the attack chain about how to gradually elicit the harmful responses from the victim model step by step. Secondly, the attacker LLM follows the attack chain to generate the initial multi-turn query set via self-talk, i.e., communicating with oneself. Finally, the attacker LLM dynamically modifies the initial attack path during the realistic interaction with the victim model. We present the concrete algorithm in Algorithm 1.
@@ 3.2 In-attack: first reason then attack
1. Infer the attack chain. Given the selected attack clue $c_{i}$, and the harmful target $x$, our attacker LLM infers a chain of thoughts $z_{1},\ldots,z_{n}$ to build the attack path from $c_{i}$ to $x$. As illustrated in Fig. 3 (a), our attack chain specifies how the topics of our multi-turn queries evolve, guiding the victim model’s responses more aligned with our attack target. In practice, each thought $z_{i}\sim p(z_{i}|x,c_{i},z_{1,\ldots,i-1};\theta)$ is sampled sequentially.

### v2

# ACL 2025 Main Conference LLMs know their vulnerabilities: Uncover Safety Gaps through Natural Distribution Shifts

@@ Abstract
Safety concerns in large language models (LLMs) have gained significant attention due to their exposure to potentially harmful data during pre-training. In this paper, we identify a new safety vulnerability in LLMs: their susceptibility to natural distribution shifts between attack prompts and original toxic prompts, where seemingly benign prompts, semantically related to harmful content, can bypass safety mechanisms. To explore this issue, we introduce a novel attack method, ActorBreaker, which identifies actors related to toxic prompts within pre-training distribution to craft multi-turn prompts that gradually lead LLMs to reveal unsafe content. ActorBreaker is grounded in Latour’s actor-network theory, encompassing both human and non-human actors to capture a broader range of vulnerabilities. Our experimental results demonstrate that ActorBreaker outperforms existing attack methods in terms of diversity, effectiveness, and efficiency across aligned LLMs. To address this vulnerability, we propose expanding safety training to cover a broader semantic space of toxic content. We thus construct a multi-turn safety dataset using ActorBreaker. Fine-tuning models on our dataset shows significant improvements in robustness, though with some trade-offs in utility. Code is available at https://github.com/AI45Lab/ActorAttack.
ACL 2025 Main Conference
LLMs know their vulnerabilities: Uncover Safety Gaps through Natural Distribution Shifts
[…]
WARNING: This paper contains model outputs that may be considered offensive.
Qibing Ren1,3  Hao Li2,3  Dongrui Liu3  Zhanxu Xie2,3  Xiaoya Lu3,1 Yu Qiao3  Lei Sha2  Junchi Yan1  Lizhuang Ma1  Jing Shao3 1MoE Key Lab of Artificial Intelligence, Shanghai Jiao Tong University 2 Beihang University    3Shanghai Artificial Intelligence Laboratory renqibing@sjtu.edu.cn  zy2442214@buaa.edu.cn  liudongrui@pjlab.org.cn
@@ 1 Introduction
Figure 1: (a): A real-world example of our multi-turn attack compared with the single-turn toxic query. (b): the schematic description of our method. Each triangle box represents an actor, semantically related to the harmful target, as a hint for our multi-turn attack. The series of white circles represent a sequence of thoughts about how to finish our multi-turn attack step by step.
Safety issues in large language models (LLMs) arise because they are pre-trained on web-scale data, which includes vast amounts of potentially harmful data. Current safety training focuses on teaching models to reject harmful queries (Ouyang et al., 2022; Bai et al., 2022; Zou et al., 2024a; Yuan et al., 2024a). However, recent research shows that these techniques can be bypassed by adversarial attacks (Zou et al., 2023b; Carlini et al., 2024) or jailbreak methods (Chao et al., 2023; Yuan et al., 2024b), by deliberately modifying harmful prompts, which we term as malicious distribution shift. This raises the robustness concern of LLMs’ safety behavior. Since LLMs have learned vast semantic relationships between harmful and neutral content during pre-training, in this paper, we ask a crucial question: how robust are aligned LLMs to natural distribution shifts, that is, benign prompts that are semantically related to harmful queries?
To answer this, we aim to identify diverse prompts from the pre-training distribution that are semantically connected to harmful content and use them to craft attacks. A motivating example is using the semantic connection between “Ted Kaczynski” (a known terrorist involved in bomb-making) and bomb-making itself. Fig. 1 (a) shows that by designing seemingly harmless queries across multiple turns, we can gradually guide the model to reveal harmful details. This example highlights a new safety vulnerability: the exploitation of benign prompts that exist within the pre-training distribution but are not covered by safety training data.
Furthermore, we introduce a multi-turn attack method, ActorBreaker, to systematically assess the robustness of aligned LLMs against these semantically related prompts. Our approach starts with a conceptual network based on Latour’s actor-network theory (Latour, 1987a), which categorizes six divergent types of actors linked to a harmful target (e.g., “Ted Kaczynski” is a node in the bomb-making network). Following Latour’s analyses, besides human actors, we also consider non-human actors like books, media, or social movements, for a better coverage of potential vulnerabilities. Fig. 1 (b) shows the overview of our method. Given a toxic query, we instantiate its actor network, including the actor’s name and its semantic relationships with the toxic query, by leveraging the pre-training prior of the LLM. Each node in the network represents a potential attack clue, and we use these clues to generate our multi-turn attack prompts. Current multi-turn attack methods either rely on specific prompt strategies, such as role-playing and scenario assumptions, or on fixed communication templates with human-designed seed instances, leading to malicious distribution shifts in their generated prompts. These methods suffer from a diversity issue due to the fixed strategies and biases in seed instances. Contrarily, our method ensures the generation of in-distribution and benign prompts without using specific jailbreak techniques.
Experimental results validate the advantages of our method in terms of diversity, effectiveness, and efficiency. ActorBreaker achieves the highest success rate on Harmbench (Mazeika et al., 2024), outperforming both leading single-turn and multi-turn attacks across aligned LLMs. Even with GPT-o1 (OpenAI, 2024b), which improves safety through advanced reasoning, our method still succeeds in generating unsafe outputs. It indicates the importance of addressing the disparity between pre-training and safety training data distributions.
Finally, to bridge this safety gap, we design adaptive defense. Rather than focusing only on specific toxic queries, we propose to expand the scope of safety alignment to cover the broader semantic space of toxic prompts. We construct a multi-turn safety alignment dataset using ActorBreaker and show that fine-tuning models on our safety dataset significantly improve their robustness against our attacks, though there is a trade-off between utility and safety. Our contributions are listed below:
1. We identify a new failure mode in aligned LLMs: their brittleness to natural distribution shifts, i.e., benign prompts that are semantically related to toxic content.
2. We propose ActorBreaker, a novel attack method for generating diverse, benign multi-turn queries related to a toxic query. Grounded in Latour’s actor-network theory, our method provides a comprehensive evaluation of LLM robustness by exploring divergent attack paths within the pre-training domain.
3. Our approach achieves state-of-the-art performance on Harmbench, outperforming both single-turn and multi-turn attack baselines. Our attack prompts bypass the detection of Llama-guard 2 (Team, 2024), demonstrating the naturalness of our prompts. Our attacks transfer well across aligned LLMs without extra optimization.
4. We demonstrate the importance of broadening the scope of safety training data to encompass the vast semantic relationships within toxic prompts. Models fine-tuned on our multi-turn safety dataset show improved robustness against our attacks.
@@ 2 Related Work
Multi-turn Attacks. Most multi-turn attack method either i) exploits specialized jailbreak techniques like hypothetical scenarios such as "The following happens in a {scenario}…" or role-playing like "You are a {role} doing {something}…"(Red queen (Jiang et al., 2024), CoA (Yang et al., 2024), CFA (Sun et al., 2024), Zhou et al. (2024c)), or ii) relies on fixed communication templates with human-designed seed instances (Crescendo (Russinovich et al., 2024), Zhou et al. (2024c)). On the one hand, fixed attack strategies and potential biases towards seed instances may lead to a diversity issue, ultimately limiting their effectiveness. On the other hand, their prompt distribution is far from ours. Their prompts are deliberately crafted using fixed strategies. But our prompts are more natural since we aim to capture the diverse semantical relationships with the toxic query in the pre-training domain. Imposter.AI (Liu et al., 2024d) utilizes question decomposition and obfuscation techniques like synonym substitution to manipulate the toxic query. Cosafe (Yu et al., 2024) proposes a multi-turn attack by using co-reference, but it proposes to directly place the harmful intent at the last query. Both methods also exploit malicious distribution shift to bypass safety mechanisms, in contrast to natural distribution shift exploited by our method. Alternatively, researchers propose to use human red teamers to manually generate multi-turn attacks using a list of human tactics (Li et al., 2024a), which is orthogonal to our work. See related work of defenses for LLMs in App. A.
3 Method
Overview. We propose a two-stage approach to automatically find attack clues and generate multi-turn attacks. The first stage consists of network construction around the seed toxic prompt, where every network node can be used as an attack clue (Fig. 2). The second stage includes the attack chain generation based on the attack clue and the multi-turn query generation (Fig. 3). We present the concrete algorithm in Algorithm 1.
@@ 3 Method
Figure 2: Druing the pre-attack stage, ActorBreaker first leverages the knowledge of LLMs to instantiate our conceptual network ${\mathcal{G}}_{concept}$ as ${\mathcal{G}}_{inst}$ as a two-layer tree. The leaf nodes of ${\mathcal{G}}_{inst}$ are specific actor names. ActorBreaker then samples actors and their relationships with the harmful target as our attack clues.
3.1 Actor-network construction
Inspired by Latour’s actor-network theory, we propose a conceptual network ${\mathcal{G}}_{concept}$ to categorize various types of actors semantically related with the seed toxic prompt and we leverage the pre-training knowledge of LLMs to specify our network.
Theoretical grounding in our actor design. Latour (1987a) claim that everything does not exist alone yet in a network of relationships, and is influenced by various actors. In the context of harmful content, different actors contribute in unique ways throughout the content lifecycle: from its creation and dissemination to its reception and regulation. As illustrated in Fig. 2, we identify six types of actors, e.g., Creation actors represent the origins of harmful ideas or inspiration and Distribution actors facilitate the spread of harmful content. We argue that the semantic relationships between these actors and the harmful prompt are encoded in the model’s knowledge and, thus can be used as our attack clues. Moreover, Latour emphasizes that human and non-human actors hold equally significant positions in the network. Therefore, for better coverage of possible attack clues, we further consider both human entities (e.g., historical figures, influential people) and nonhuman entities (e.g., books, media, social movements) within each category of actors. Our categorization is consistent with other applications of ANT Callon (1984); Latour (1987b).
Network Definition. Our network is a two-layered tree structure, where the root node is the harmful target $x$. First layer consists of six abstract types of actors. Leaf nodes are specific actor names. Each edge captures the semantic relationship between an actor and the harmful target, which forms a potential attack clue $c_{i}$.
Network adaptation to new harmful targets. We generate a unique network for each harmful target, ensuring the derived clues are semantically relevant to the given target. As illustrated in Figure 2, we instruct LLMs to automatically instantiate nodes and edges of the network as ${\mathcal{G}}_{inst}$, based on our conceptual descriptions of the network ${\mathcal{G}}_{concept}$ and the harmful target $x$, that is, ${\mathcal{G}}_{inst}\sim p(x,{\mathcal{G}}_{concept};\theta)$. Finally, we extract our diverse attack clue set $C$=$[c_{1},\ldots,c_{n}]$ from ${\mathcal{G}}_{inst}$, that is, $C\sim{\mathcal{G}}_{inst}$.
3.2 In-attack
Based on the constructed network, we perform our multi-turn attacks in three steps. The first step is to infer the attack chain about how to gradually elicit the harmful responses from the victim model step by step. Secondly, the attacker LLM follows the attack chain to generate the initial multi-turn query set via self-talk, i.e., communicating with oneself. Finally (optional), the attacker LLM dynamically modifies the initial attack path during the realistic interaction with the victim model.
@@ 3.2 In-attack
1. Infer the attack chain. Given the selected attack clue $c_{i}$ and the harmful target $x$, our attacker LLM infers a chain of thoughts $z_{1},\ldots,z_{n}$ to build the attack path from $c_{i}$ to $x$. As Fig. 3 (a) shows, our attack chain specifies how the topics of our multi-turn queries evolve, guiding the victim model’s responses more aligned with our attack target. In practice, each thought $z_{i}\sim p(z_{i}|x,c_{i},z_{1,\ldots,i-1};\theta)$ is sampled sequentially.